OpenStudy (anonymous):

Find the unit vector in the direction of u, and write your answer in component form. Let u = <-8, -9>. Find 7u.

OpenStudy (anonymous):

Please help !!

OpenStudy (anonymous):

@amistre64 please help

OpenStudy (anonymous):

@phi @ganeshie8 please help

OpenStudy (zzr0ck3r):

do you know how to find the magnitude of u?

OpenStudy (zzr0ck3r):

@Andreshoyos23 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but i belive i figured it out @zzr0ck3r

OpenStudy (zzr0ck3r):

if you have a vector (a,b) then the magnitude is \[\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

is it <-56, -63>

OpenStudy (zzr0ck3r):

that is the answer to the second part

OpenStudy (zzr0ck3r):

but the first part wants you to find the unit vector

OpenStudy (anonymous):

the magnitude is squared and rooted?

OpenStudy (anonymous):

meaning if i have the two numbers -8 and -9 you must square and muliply by two

OpenStudy (zzr0ck3r):

\[given \space <a,b>\\\text{then the unit vector is}\\\frac{1}{\sqrt{a^2+b^2}}*<a,b>=<\frac{a}{\sqrt{a^2+b^2}},\frac{b}{\sqrt{a^2+b^2}}>\]

OpenStudy (anonymous):

ah !!!! ok :D i see